OpenStudy (anonymous):

Find the condition that the equation 1/x+1/(x+a)=1/m+1/(m+a) may have roots equal in magnitude but opposite in sign.

OpenStudy (anonymous):

\[a=m \sqrt{2}\]

OpenStudy (anonymous):

\[ax^2+bx+c=0\]

OpenStudy (anonymous):

STEP BY STEP PLEASE

OpenStudy (anonymous):

Sory man here the answer is 2m^2=a^2

OpenStudy (anonymous):

same thing..

OpenStudy (anonymous):